OpenStudy (anonymous):

If the area of an isosceles right triangle is 36, what is its perimeter?

OpenStudy (angela210793):

A=x^2/2=36 x^2=72 x=sqr72=6sqr2 hypotenuse will be 12(Pythagorean theorem) Perimeter=12sqr2+12=12(sqr2+1)

OpenStudy (anonymous):

Thanx

OpenStudy (angela210793):

uw :)